# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ III-D Switching Law
Assume that the optimal trajectory in problem (1a) consists of a finite number of arcs represented as ${\mathcal{S}}_{1},{\mathcal{S}}_{2},\dots,{\mathcal{S}}_{N}$ successively. The switching law of the optimal trajectory is defined as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$.
Remark.
The switching law focuses on the arcs that compose the optimal trajectory, while the motion time of each arc is not included in a switching law. Furthermore, Definition 2 does not characterize the chattering phenomenon since chattering induces an infinite number of arcs.
The connection between arcs is analyzed in Proposition 3. Denote the two adjacent arcs as ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$, where ${\mathcal{S}}_{1}\not={\mathcal{S}}_{2}$. $\forall i=1,2$, ${\mathcal{S}}_{i}=\left(\widehat{{\boldsymbol{A}}}_{i},\widehat{{\boldsymbol{b}}}_{i},\widehat{{\boldsymbol{F}}}_{i},\widehat{{\boldsymbol{g}}}_{i},\mathcal{P}_{i}\right)$. Among them, ${\mathcal{S}}_{1}={\mathcal{S}}_{2}$ is defined as $\widehat{{\boldsymbol{A}}}_{1}=\widehat{{\boldsymbol{A}}}_{2}$, $\widehat{{\boldsymbol{b}}}_{1}=\widehat{{\boldsymbol{b}}}_{2}$, and $\mathcal{P}_{1}=\mathcal{P}_{2}$ since $\widehat{{\boldsymbol{F}}}_{i}$ and $\widehat{{\boldsymbol{g}}}_{i}$ can be induced by $\mathcal{P}_{i}$. Before introducing Proposition 3, a lemma is provided as preparation.
Lemma 2.
Consider the IVP $\dot{\boldsymbol{x}}\left(t\right)={\boldsymbol{A}}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{b}}$, $t\in{\mathbb{R}}$, where ${\boldsymbol{x}}\left(t_{0}\right)={\boldsymbol{x}}_{0}\in{\mathbb{R}}^{n}$, ${\boldsymbol{A}}\in{\mathbb{R}}^{n\times n}$, ${\boldsymbol{b}}\in{\mathbb{R}}^{n}$. If ${\boldsymbol{c}}\in{\mathbb{R}}^{n}$, s.t. $\forall 1\leq r\leq n$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\mathrm{const}}$ on ${\mathbb{R}}$.
Proof.
According to Lemma 1, $\forall t\in{\mathbb{R}}$, ${\boldsymbol{x}}\left(t\right)={\mathrm{e}}^{{\boldsymbol{A}}\left(t-t_{0}\right)}{\boldsymbol{x}}_{0}+\left(\int_{t_{0}}^{t}{\mathrm{e}}^{{\boldsymbol{A}}\left(t-\tau\right)}\left(\tau\right)\mathrm{d}\tau\right){\boldsymbol{b}}$. Considering the Jordan normal form of ${\boldsymbol{A}}$, it has $\forall t\in{\mathbb{R}}$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{i=1}^{l}p_{i}\left(t-t_{0}\right){\boldsymbol{e}}^{\lambda_{i}\left(t-t_{0}\right)}$, where $\left\{\lambda_{i}\right\}_{i=1}^{l}$ are the eigenvalues of ${\boldsymbol{A}}$, and $p_{i}$ is a polynomial of degree not greater than $n$. Evidently, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}$ is an analytic function; hence, $\forall t\in{\mathbb{R}}$, it has
${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)=\sum_{r=0}^{\infty}\frac{\left(t-t_{0}\right)^{r}}{r!}\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)\right|_{t_{0}}.$ (18)
The case where ${\boldsymbol{A}}={\boldsymbol{0}}$ is trivial. Assume that ${\boldsymbol{A}}\not={\boldsymbol{0}}$. Considering the minimal polynomial of ${\boldsymbol{A}}$, $\exists p^{*}$, s.t. $p^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{n}=p^{*}\left({\boldsymbol{A}}\right)$. Therefore, $\forall r\geq n$, $\exists p_{n}^{*}$ s.t. $p_{n}^{*}$ is a polynomial of degree less than $n$ and ${\boldsymbol{A}}^{r}=p_{n}^{*}\left({\boldsymbol{A}}\right)$.
Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}+{\boldsymbol{b}}\right)$. Since $\forall 1\leq r\leq n$, $\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, it holds that $\forall r\geq n+1$,
$\frac{\mathrm{d}^{r}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)={\boldsymbol{c}}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)$ (19)
$=$ ${\boldsymbol{c}}^{\top}p_{r-1}^{*}\left({\boldsymbol{A}}\right)\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{0}\right)+{\boldsymbol{b}}\right)=0.$
By (18), $\forall t\in{\mathbb{R}}$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)\equiv{\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{0}\right)$. ∎
@@
Assume that $\forall i=1,2$, ${\mathcal{S}}_{i}$ occurs for $t\in\left[t_{i-1},t_{i}\right]$ in the above notation. ${\boldsymbol{x}}$ enters ${\mathcal{S}}_{2}$ from ${\mathcal{S}}_{1}$ at $t_{1}$.
@@
    If ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ are constrained arcs, then $\mathcal{P}_{1}\cap\mathcal{P}_{2}=\varnothing$. Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (20)
    $\forall p\in\mathcal{P}_{2}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (21)
@@
    If ${\mathcal{S}}_{1}$ is a constrained arc and ${\mathcal{S}}_{2}$ is an unconstrained arc, then $\forall p,q\in\mathcal{P}_{1}$, ${\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right)={\mathrm{sgn}}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{A}}^{r_{q}-1}{\boldsymbol{b}}\right)$. On ${\mathcal{S}}_{2}$, $\forall p\in\mathcal{P}_{1}$, it has
@@
    Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)=0.\end{dcases}$ (23)
@@
    If ${\mathcal{S}}_{1}$ is an unconstrained arc and ${\mathcal{S}}_{2}$ is a constrained arc, then $\forall p,q\in\mathcal{P}_{2}$, ${\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right)={\mathrm{sgn}}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{A}}^{r_{q}-1}{\boldsymbol{b}}\right)$. On ${\mathcal{S}}_{1}$, $\forall p\in\mathcal{P}_{2}$, it has
@@
    Furthermore, $\forall p\in\mathcal{P}_{2}$, $\exists r_{p}\leq\hat{r}_{p}\leq n$, s.t.
    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)<0,\\ \forall 1\leq r<\hat{r}_{p},\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)=0.\end{dcases}$ (25)
@@
Consider the connection between two unconstrained arcs. By Proposition 1, assume that $\forall i=1,2$, $u\equiv u_{i}$ on ${\mathcal{S}}_{i}$, where $u_{i}\in\left\{\pm{u_{\mathrm{m}}}\right\}$. Then, $\widehat{{\boldsymbol{b}}}_{i}=u_{i}{\boldsymbol{b}}$. Since $\widehat{{\boldsymbol{A}}}_{1}=\widehat{{\boldsymbol{A}}}_{2}={\boldsymbol{A}}$ and $\mathcal{P}_{1}=\mathcal{P}_{2}=\varnothing$, it holds that $\widehat{{\boldsymbol{b}}}_{1}\not=\widehat{{\boldsymbol{b}}}_{2}$; hence, $u_{1}=-u_{2}$. Therefore, Proposition 3-1 holds.
Consider the connection between two constrained arcs. Assume that $\mathcal{P}_{1}\cap\mathcal{P}_{2}\not=\varnothing$. Arbitrarily take a $p^{*}\in\mathcal{P}_{1}\cap\mathcal{P}_{2}$ into consideration. According to Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}-\frac{{\boldsymbol{b}}{\boldsymbol{c}}_{p^{*}}^{\top}{\boldsymbol{A}}^{r_{p^{*}}}}{{\boldsymbol{c}}_{p^{*}}^{\top}{\boldsymbol{A}}^{r_{p^{*}}-1}{\boldsymbol{b}}}\right){\boldsymbol{x}}$ holds. Evidently, $\forall p\in\mathcal{P}_{1}\cup\mathcal{P}_{2}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}\equiv 0$, and $\forall 1\leq r<r_{p}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)\equiv 0$. By Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds. Therefore, $\mathcal{P}_{1}=\mathcal{P}_{2}$, which contradicts ${\mathcal{S}}_{1}\not={\mathcal{S}}_{2}$. Hence, $\mathcal{P}_{1}\cap\mathcal{P}_{2}=\varnothing$.
Furthermore, consider the monotonicity of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+{\boldsymbol{d}}_{p}$ at $t_{1}$ to guarantee the feasibility of ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$. $\exists\varepsilon>0$, s.t. $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{d}}_{p}<0$. Note that on ${\mathcal{S}}_{2}$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{2}{\boldsymbol{x}}$ and $\forall r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{r}{\boldsymbol{x}}\left(t_{1}\right)$. Lemma 2 implies that $\exists 1\leq\hat{r}_{p}\leq n$, s.t. $\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall 1\leq r<\hat{r}_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$. Since $\forall 1\leq r<r_{p}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, it has $\hat{r}_{p}\geq r_{p}$. In other words, (20) holds. For the same reason, (21) holds.
@@
$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall 1\leq r<\hat{r}_{p},\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (26)
Hence, $\exists\varepsilon>0$, $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$. For the same reason, $\exists\varepsilon^{\prime}\in\left(0,\varepsilon\right)$, $\forall p\in\mathcal{P}_{2}$, $t\in\left(t_{1}-\varepsilon^{\prime},t_{1}\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$. Hence, Proposition 3-2 holds.

### v2

@@ III-D Switching Law
Assume that a BBS trajectory in problem (1a) consists of a finite number of arcs ${\mathcal{S}}_{1},{\mathcal{S}}_{2},\dots,{\mathcal{S}}_{N}$ successively. The switching law is denoted as ${\mathcal{L}}={\mathcal{S}}_{1}{\mathcal{S}}_{2}\dots\mathcal{S}_{N}$.
The switching law focuses on the arcs that compose the optimal trajectory, while the motion time of each arc is not included in a switching law. For the example in Fig. 1, the switching law is ${\mathcal{S}}_{1}{\mathcal{S}}_{2}{\mathcal{S}}_{3}$. ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{3}$ are unconstrained arcs where ${\boldsymbol{x}}>{\boldsymbol{0}}$ holds a.e. ${\mathcal{S}}_{2}$ is a constrained arc where $x_{2}\equiv 0$.
The following proposition discusses connections of two adjacent arcs, including unconstrained and constrained arcs.
@@
Assume that ${\mathcal{S}}_{i}=\left(\widehat{{\boldsymbol{A}}}_{i},\widehat{{\boldsymbol{b}}}_{i},{\boldsymbol{F}}_{i},{\boldsymbol{g}}_{i},\mathcal{P}_{i}\right)$ occurs in $\left[t_{i-1},t_{i}\right]$, $i=1,2$, where $\left[\widehat{{\boldsymbol{A}}}_{1},\widehat{{\boldsymbol{b}}}_{1}\right]\not=\left[\widehat{{\boldsymbol{A}}}_{2},\widehat{{\boldsymbol{b}}}_{2}\right]$. The following conclusions hold at the connection of ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$.
@@
    If ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$ are constrained arcs, then $\mathcal{P}_{1}\cap\mathcal{P}_{2}=\varnothing$. Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t.
    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (19)
    $\forall p\in\mathcal{P}_{2}$, $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t.
    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{\hat{r}_{p}}{\boldsymbol{x}}\left(t_{1}\right)<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{1}^{r}{\boldsymbol{x}}\left(t_{1}\right)=0.\end{dcases}$ (20)
@@
    If ${\mathcal{S}}_{1}$ is a constrained arc and ${\mathcal{S}}_{2}$ is an unconstrained arc, then $\forall p,q\in\mathcal{P}_{1}$, ${\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right)={\mathrm{sgn}}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{A}}^{r_{q}-1}{\boldsymbol{b}}\right)$. On ${\mathcal{S}}_{2}$, $\forall p\in\mathcal{P}_{1}$, it holds that
@@
    Furthermore, $\forall p\in\mathcal{P}_{1}$, $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t.
    $\begin{dcases}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{2}{\boldsymbol{b}}\right)=0.\end{dcases}$ (22)
@@
    If ${\mathcal{S}}_{1}$ is an unconstrained arc and ${\mathcal{S}}_{2}$ is a constrained arc, then $\forall p,q\in\mathcal{P}_{2}$, ${\mathrm{sgn}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}\right)={\mathrm{sgn}}\left({\boldsymbol{c}}_{q}^{\top}{\boldsymbol{A}}^{r_{q}-1}{\boldsymbol{b}}\right)$. On ${\mathcal{S}}_{1}$, $\forall p\in\mathcal{P}_{2}$, it holds that
@@
    Furthermore, $\forall p\in\mathcal{P}_{2}$, $\exists\hat{r}_{p}\in\left[r_{p},n\right]\cap{\mathbb{N}}$, s.t.
    $\begin{dcases}\left(-1\right)^{\hat{r}_{p}}{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{\hat{r}_{p}-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r-1}\left({\boldsymbol{A}}{\boldsymbol{x}}\left(t_{1}\right)+u_{1}{\boldsymbol{b}}\right)=0.\end{dcases}$ (24)
@@
Consider the connection between two unconstrained arcs. By Proposition 1, assume that $\forall i=1,2$, $u\equiv u_{i}$ on ${\mathcal{S}}_{i}$, where $u_{i}\in\left\{\pm{u_{\mathrm{m}}}\right\}$. Then, $\widehat{{\boldsymbol{b}}}_{i}=u_{i}{\boldsymbol{b}}$. Since $\widehat{{\boldsymbol{A}}}_{1}=\widehat{{\boldsymbol{A}}}_{2}={\boldsymbol{A}}$ and $\mathcal{P}_{1}=\mathcal{P}_{2}=\varnothing$, it holds that $\widehat{{\boldsymbol{b}}}_{1}\not=\widehat{{\boldsymbol{b}}}_{2}$; hence, $u_{1}=-u_{2}$. Therefore, Proposition 4.1 holds.
Consider the connection between two constrained arcs. Assume that $\mathcal{P}_{1}\cap\mathcal{P}_{2}\not=\varnothing$. Arbitrarily take a $p^{*}\in\mathcal{P}_{1}\cap\mathcal{P}_{2}$ into consideration. According to Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}-{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p^{*}}^{\top}\right){\boldsymbol{x}}$ holds. Evidently, $\forall p\in\mathcal{P}_{1}\cup\mathcal{P}_{2}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}\equiv 0$, and $\forall r\in\left[r_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)\equiv 0$. By Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds. Therefore, $\mathcal{P}_{1}=\mathcal{P}_{2}$, which contradicts ${\mathcal{S}}_{1}\not={\mathcal{S}}_{2}$. Hence, $\mathcal{P}_{1}\cap\mathcal{P}_{2}=\varnothing$.
Furthermore, consider the monotonicity of ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+{\boldsymbol{d}}_{p}$ at $t_{1}$ to guarantee the feasibility of ${\mathcal{S}}_{1}$ and ${\mathcal{S}}_{2}$. $\exists\varepsilon>0$, s.t. $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+{\boldsymbol{d}}_{p}<0$. Note that on ${\mathcal{S}}_{2}$, $\dot{\boldsymbol{x}}=\widehat{{\boldsymbol{A}}}_{2}{\boldsymbol{x}}$ and $\forall r\in{\mathbb{N}}^{*}$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}={\boldsymbol{c}}_{p}^{\top}\widehat{{\boldsymbol{A}}}_{2}^{r}{\boldsymbol{x}}\left(t_{1}\right)$. Lemma 1 implies that $\exists\hat{r}_{p}\in\left[n\right]$, s.t. $\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0$ and $\forall r\in\left[\hat{r}_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$. Since $\forall r\in\left[r_{p}-1\right]$, $\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0$, it holds that $\hat{r}_{p}\geq r_{p}$. In other words, (19) holds. For the same reason, (20) holds.
@@
$\begin{dcases}\left.\frac{\mathrm{d}^{\hat{r}_{p}}}{\mathrm{d}t^{\hat{r}_{p}}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}<0,\\ \forall r\in\left[\hat{r}_{p}-1\right],\,\left.\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\right)\right|_{t_{1}^{+}}=0.\end{dcases}$ (25)
Hence, $\exists\varepsilon>0$, $\forall p\in\mathcal{P}_{1}$, $t\in\left(t_{1},t_{1}+\varepsilon\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$. For the same reason, $\exists\varepsilon^{\prime}\in\left(0,\varepsilon\right)$, $\forall p\in\mathcal{P}_{2}$, $t\in\left(t_{1}-\varepsilon^{\prime},t_{1}\right)$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$. Hence, Proposition 4.2 holds.